Two planes take off from Salt Lake City, UT. The first plane is traveling 2.5 times faster than the second plane. After traveling in the same direction for 9.5 hours, they are 1239.75 miles apart. what is the average speed of each plane?
Answer by checkley77(12844)

2.5x-x=1239.75
1.5x=1239.75
x=1239.75/1.5
x=826.5 mph. for the slower plane.
2.5*826.5=2,066.25 for the faster plane.
Proof:
2,055.25-826.50=1,239.75